At a tennis tournament a statistician kept track of every serve. The statistician reported that the mean serve speed of a particular player was 97 miles per hour (mph) and the standard deviation of the serve speeds was 11 mph. Assume that the statistician also gave us the information that the distribution of the serve speeds was bell shaped. What proportion of the playerʹs serves are expected to be between 119 mph and 130 mph?

The proportion is ____ (round to 4 decimal places as needed)
Answer #1

ANSWER:

Given that,

At a tennis tournament a statistician kept track of every serve. The statistician reported that the mean serve speed of a particular player was 97 miles per hour (mph) and the standard deviation of the serve speeds was 11 mph. Assume that the statistician also gave us the information that the distribution of the serve speeds was bell shaped. What proportion of the playerʹs serves are expected to be between 119 mph and 130 mph?

Mean = \mu = 97 mph

Standard deviation = \sigma = 11 mph

z = (X-\mu)/\sigma = (X-97)/11

P(119 < X < 130) = P((119-97)/11 < (X-\mu)/\sigma < (130-97)/11)

P(119 < X < 130) = P(22/11 < Z < 33/11)

P(119 < X < 130) = P(2 < Z < 3)

P(119 < X < 130) = P(Z < 3)-P(Z < 2)

P(119 < X < 130) = 0.99865-0.97725 (From z score table)

P(119 < X < 130) = 0.0214

0.1 1.3 STANDARD NORMAL DISTRIBUTION: Table Values Represent AREA to the LEFT of the Z score. Z .00 .01 .02 .03 .04 .05 .06 .